OpenStudy (anonymous):

Please help me multipy (4x+3y)^2

OpenStudy (anonymous):

First we need to find out what x equals or y equals...it looks like it will be easy to get y alone in the second equation: 4x=3y-2 18=3x+y 18-3x=y Now that we know y equals 18-3x I can plug in 18-3x anywhere I see a y 4x=3(18-3x)-2 Distribute the 3 4x=54-9x-2 combine like terms 4x=52-9x Add 9x to both sides 13x=52 divide both sides by 13 x=4 Now that I know x=4 I can plug it into one of the equations: 4*4=3y-2 16=3y-2 Add 2 to both sides 18=3y divide both sides by 3 y=6

OpenStudy (anonymous):

medal plz

OpenStudy (anonymous):

MAli13chineta, thanks, and how do I give medals?

OpenStudy (anonymous):

on my comment on the right click the little medal above the word delete